# Purified pseudomode model for nonlinear system-bath interactions

### v1

@@ Abstract
The theory of purified pseudomodes [arXiv:2412.04264 (2024)] was recently developed to provide a numerical tool for the analysis of the properties of a quantum system and the environment it couples to via linear system-bath interactions. Here we extend this theory to allow for the description of general nonlinear system-bath interactions. We demonstrate the validity of our method by considering the spontaneous decay of a two-level atom placed inside a single-mode lossy cavity and furthermore, its potential application to nanophotonics by calculating the resonance fluorescence spectrum of a quantum dot in the presence of a phonon environment. Our method provides a useful tool for the study of phonon-assisted emission in quantum dots and holds the the promise for broad applications in fields like quantum biology, nonlinear phononics, and nanophotonics.
@@
The manipulation and synthesis of photon- and phonon-matter interactions have achieved considerable progresses in past decades, making some challenging regimes experimentally accessible. Representative examples include, e.g., ultra- and deep-strong light-matter interactions in cavity and circuit quantum electrodynamics (QED) [1, 2, 3], time retardation effects caused by light propagation in waveguide QED [4, 5, 6] and giant atoms [7, 8, 9, 10], nonlinear optomechanical interactions induced by radiation pressure [11, 12, 13, 14], and nonlinear phononics [15, 16, 17], to name a few. A faithful and accurate characterization of the properties of physical systems in said regimes requires the improvement of existing or the invention of new analytical and numerical methods [18, 19].
@@
The common starting point of devising such methods is often a microscopic description of the environment and its interaction with the system of interest, and a variety of perturbative or non-perturbative treatments can be derived depending on the parameter regime of interest. For example, perturbation theory can lead to an effective description of the system alone in terms of quantum master equations [18, 19, 20, 21, 22] in which the environmental effects are implicitly accounted for. On the other hand, non-perturbative methods can be grounded on a variety of ideas, ranging from discretization of the the continous environment [23, 24, 25] or the Feymann-Vernon influence functional [26, 27, 28, 29, 30, 31, 32, 33] for tensor-network simulations to solving unitarily equivalent models obtained from the reaction coordinate mapping [34, 35, 36, 37, 38, 39], the chain mapping [40, 41, 42, 43] and the polaron transformation [44, 45, 46, 47, 48, 49, 50], or constructing effective models with auxiliary degrees of freedom, e.g., the hierarchical equations of motion (HEOM) [51, 52, 53, 54, 55, 56, 57, 58, 59, 60], dissipatons [61, 62, 63, 64], and pseudomodes [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78].
@@
In a previous paper [79], based on the pseudomode theory [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78], some of us introduced a new class of ancillary damping modes and detailed a rigorous scheme to construct effective models allowing for practical simulations of the dynamics of a system and the input-output properties of its environment. These ancillary modes were named “purified pseudomodes” because their states effectively remain pure during the evolution, in stark contrast to pseudomodes which require a mixed-state description. This method is numerically exact and valid for thermal bosonic environments and for generic linear system-bath interactions, and only requires the spectral decompositions of the bath two-time correlation matrix.
In this paper, we aim to derive an extended purified pseudomode model for the simulation of open quantum systems in the presence of general nonlinear system-bath interactions, where the bath coupling operator is an analytical function of a linear coupling operator $X$, see the definition in Eq. (2). This topic remains largely unexplored in the literature, except for Refs. [80, 81] where the specific case of a system-bath interaction written as the sum of a linear term and a quadratic correction was considered in the context of the HEOM.
@@
The analysis of nonlinear system-bath interactions, including the specific ones employed in Refs. [80, 81], is more challenging than the linear case because of the additional terms emerging in the Wick factorizations of multi-time bath correlation functions which depend on non-trivial equal-time contractions. Here, we show that the technical difficulties involved in re-summing all these extra Wick contractions due to nonlinearities are bypassed using the pseudomode theory. In fact, by its very definition, this method offers a natural one-to-one correspondence of multi-time correlation functions between the original bath and the corresponding pseudomode model, making the latter able to encode any nonlinear statistics of the original bath by mimicking the statistical properties on the linear operator $X$. We provide an easily implementable scheme to define the required pseudomode model and show that a further purification of the degrees of freedom of this model allows to optimize the overall memory requirements. The validity of our method is demonstrated for the problem of the spontaneous decay of a two-level atom strongly coupled to a single-mode lossy cavity. We also illustrate the potential applications of our method to more complex examples by studying the effects of nonlinearly-coupled phonons on the fluorescence spectrum of a two-level quantum dot.
@@
where $H_{S}$ is the system Hamiltonian and $H_{B}=\sum_{k}\omega_{k}b_{k}^{\dagger}b_{k}$ is the bath Hamiltonian with $b_{k}$ ($b_{k}^{\dagger}$) denoting the annihilation (creation) operator of the bath harmonic mode having frequency $\omega_{k}$, satisfying $[b_{k},b_{k^{\prime}}^{\dagger}]=\delta_{kk^{\prime}}$. We consider the following general form for the system-bath interaction
@@
where $\rho_{S}^{I}(t)=e^{iH_{0}t}\rho_{S}(t)e^{-iH_{0}t}$ and $H_{\text{int}}^{I}(t)=e^{iH_{0}t}H_{\text{int}}e^{-iH_{0}t}$ denote the corresponding reduced density operator and system-bath interaction Hamiltonian in the interaction picture, respectively. The formal solution of Eq. (3), namely, the Dyson series can be written as [73]
@@
with $\langle\mathcal{Q}_{t_{m}}^{i_{m}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}\equiv\text{Tr}_{B}\Big{[}\mathcal{T}_{B}\mathcal{Q}_{t_{m}}^{i_{m}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rho_{\text{th}}\Big{]}$ denoting the multi-time bath correlation functions and the indexes $i_{1},\cdots,i_{m}\in\{l,r\}$ labelling the left- or right-action of the superoperators $\mathcal{S}_{i_{j}}^{t_{j}}$ and $\mathcal{Q}_{t_{j}}^{i_{j}}$. In other words, they are defined as $\mathcal{S}_{t_{j}}^{l}[\cdot]=s(t_{j})[\cdot]$, $\mathcal{S}_{t_{j}}^{r}[\cdot]=-[\cdot]s(t_{j})$, $\mathcal{Q}_{t_{j}}^{l}[\cdot]=Q(t_{j})[\cdot]$, and $\mathcal{Q}_{t_{j}}^{r}[\cdot]=[\cdot]Q(t_{j})$ where $s(t_{j})=\exp(iH_{B}t_{j})s\exp(-iH_{B}t_{j})$ and $Q(t_{j})=\exp(iH_{B}t_{j})Q\exp(-iH_{B}t_{j})$ correspond to the operators $s$ and $Q$ in the interaction picture, respectively. Note that, the minus sign in the definition of $\mathcal{S}_{t_{j}}^{r}$ is necessary to recover the commutators involving $H_{\text{int}}^{I}(t)$ in the Dyson series in Eq. (4). The time-ordering operators $\mathcal{T}_{S}$ and $\mathcal{T}_{B}$ ensure that the superoperators $\mathcal{S}_{i_{j}}^{t_{j}}$ and $\mathcal{Q}_{t_{j}}^{i_{j}}$ are in chronological order, respectively.
[…]
where $t=t_{2}-t_{1}$ denote the time difference and $J(\omega)\equiv\pi\sum_{k}\lvert g_{k}\rvert^{2}\delta(\omega-\omega_{k})$ is the spectral density. The time-translational invariance of $C(t)$ is a consequence of the stationarity of the bath initial state $\rho_{\text{th}}$ under the bath free dynamics, manifested in the commutation relation $[H_{B},\rho_{\text{th}}]=0$. In addition, $C(t)$ satisfies the time-reversal property $C(t)=C^{*}(-t)$ with the symbol * representing the complex conjugation. Ultimately, the formal series in Eq. (4) can be re-summed to yield a compact expression $\rho_{S}^{I}(t)=\exp(-\mathcal{F}(t,s,C(t))\rho_{S}(0)$ where the influence superoperator $\mathcal{F}(t,s,C(t)$ is explicitly written as [73]
$\begin{array}[]{lll}\mathcal{F}[\boldsymbol{\cdot}]&\equiv&\displaystyle\int_{0}^{t}dt_{2}\int_{0}^{t_{2}}dt_{1}\Big{(}C(t_{1}-t_{2})[\boldsymbol{\cdot}s(t_{1}),s(t_{2})]\\ &&\displaystyle~{}~{}~{}-C(t_{2}-t_{1})[s(t_{1})\boldsymbol{\cdot},s(t_{2})]\Big{)}.\end{array}$ (7)
@@
While no compact expression can be found for $\rho_{S}^{I}(t)$ in the general case of nonlinear system-bath interactions considered in Eq. (2), Wick’s theorem [73] still allows to conclude that each multi-time correlation function $\langle\mathcal{Q}_{t_{n}}^{i_{n}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{th}$ in Eq. (5) can be written as a sum of total contraction of the linear operator $X$. In this case, the resulting factorizations of $\langle\mathcal{Q}_{t_{n}}^{i_{n}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}$ acquire explicit dependencies on the second moment of $X$, i.e.
@@
As an example, the factorization of $\langle\mathcal{Q}_{t_{2}}^{i_{2}}\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}$ reads
@@
where the moment $\sigma^{2}$ and the two-time correlation function $C(t)$ appear in the first and second lines, respectively. This explicit dependence on the moment $\sigma^{2}$ in the Wick factorizations of $\langle\mathcal{Q}_{t_{n}}^{i_{n}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}$, has critical and profound consequences on the construction of the purified pseudomode model, constituting a major challenge which we overcome in the next section.
@@ III Purified pseudomode model for nonlinear system-bath interactions
In this section, we explicitly construct a purified pseudomode model that allows to reproduce the reduced system dynamics of the nonlinear open-system model described in Sec. II. To this end, we first present a brief summary of the conventional pseudomode theory in Sec. III.1. We then establish an explicit pseudomode model that reproduces both the two-time correlation (6) and the second moment (8) in Sec. III.2, based on spectral decompositions of the two-time correlation (6). In Sec. III.3, we generalize the purification technique in our previous work [79] to reach a purified pseudomode model for the special case $Q=X^{n}$ and ultimately, in Sec. III.4, we derive the desired purified pseudomode model for the general nonlinear system-bath interaction (2).
@@ III.1 The pseudomode theory
and $d_{\alpha}$ ($d_{\alpha}^{\dagger}$) denote the annihilation (creation) operator of the $\alpha$th pseudomode with frequency $\Omega_{\alpha}$ and decay rate $\Gamma_{\alpha}$. In Eq. (10), we have used a simplified version of the general pseudomode model [68, 69, 73], namely, each pseudomode is in contact with its own background environment held at zero temperature, resulting in a single Lindblad dissipator with decay rate $\Gamma_{\alpha}$ in the second line of Eq. (10). For consistency, this also implies that the initial state of the pseudomode model in Eq. (10) should be $\rho_{\text{pm}}(0)=\rho_{S}\otimes\rho_{\text{vac}}$ with $\rho_{\text{vac}}=\otimes_{\alpha}|0_{\alpha}\rangle\langle 0_{\alpha}|$ the vacuum state of all pseudomodes. Note that this zero-temperature assumption, alongside the independent pseudomode assumption underlying Eq. (10), can both be relaxed to further enlarge predictivity of the pseudomode model [69, 70, 73, 72, 74]. For example, all model parameters can be analytically continued to the complex plane such that they become complex-valued, i.e., $\Omega_{\alpha},\Gamma_{\alpha},\lambda_{\alpha}\in\mathbb{C}$ [69, 73]. On the other hand, mode-mode interactions can also be introduced to form a pseudomode network [70, 72, 74].
@@
where $C_{\text{pm}}(t)\equiv\text{Tr}_{\text{pm}}[X_{\text{pm}}(t)X_{\text{pm}}\rho_{\text{vac}}]$ can be solved analytically as [73]
@@
The above consideration can be easily generalized to nonlinear system-bath interactions (2). This is achieved by simply substituting in Eq. (10)
$Q_{\text{pm}}=\sum_{n=1}^{\infty}\alpha_{n}X_{\text{pm}}^{n},$ (14)
and again by satisfying the equivalence condition (12).
It is important to note that the equivalence in Eq. (12) is not required to be point-like. In fact, $C(t)$ always appears in the integrand of the integral in Eq. (5), meaning that the difference $\lvert C(t)-C_{\text{pm}}(t)\rvert$ can even be allowed to be arbitrarily large over a countable collection of times $\mathcal{S}=\{t|\lvert C(t)-C_{\text{pm}}(t)\rvert~{}\text{can be arbitrarily large}\}$. From this formal point of view, the explicit emergence of the moment in Eq. (8) in the Wick factorizations of the multi-time correlation function $\langle\mathcal{Q}_{t_{m}}^{i_{m}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}$ for the nonlinear interaction (2) requires a precise approximation of the correlation at $t=0$, i.e., $(t=0)\notin\mathcal{S}$, in stark constrast to the linear case. In Sec. III.2 we show that this crucial distinction leads to an extra zero-frequency pseudomode with respect to the linear case.

### v2

@@ Abstract
The theory of purified pseudomodes [arXiv:2412.04264 (2024)] was recently developed to provide a numerical tool for the analysis of the properties of a quantum system and the environment it couples to via linear system-bath interactions. Here we extend this theory to allow for the description of general nonlinear system-bath interactions. We demonstrate the validity of our method by considering the spontaneous decay of a two-level atom placed inside a single-mode lossy cavity and furthermore, its potential application to nanophotonics by calculating the resonance fluorescence spectrum of a quantum dot in the presence of a phonon environment. Our method provides a useful tool for the study of phonon-assisted emission in quantum dots and holds the promise for broad applications in fields like quantum biology, nonlinear phononics, and nanophotonics.
@@
The manipulation and synthesis of photon- and phonon-matter interactions have achieved considerable progresses in past decades, making some challenging regimes experimentally accessible. Representative examples include, e.g., ultra- and deep-strong light-matter interactions in cavity and circuit quantum electrodynamics (QED) [1, 2, 3], time retardation effects caused by light propagation in waveguide QED [4, 5, 6] and giant atoms [7, 8, 9, 10], nonlinear optomechanical interactions induced by radiation pressure [11, 12, 13, 14], and nonlinear phononics [15, 16, 17], to name a few. A faithful and accurate characterization of the properties of physical systems in these regimes requires specific analytical and numerical methods.
@@
The common starting point of devising such methods is often a microscopic description of the environment and its interaction with the system of interest [18, 19]. A variety of perturbative or non-perturbative treatments can be derived depending on the parameter regime of interest. For example, perturbation theory can lead to an effective description of the system in terms of quantum master equations [18, 19, 20, 21, 22]. On the other hand, to model non-pertubative regimes, more advanced techniques has to be used such as the discretization of the environmental continuum  [23, 24, 25] or the Feymann-Vernon influence functional [26, 27, 28, 29, 30, 31, 32, 33] for tensor-network simulations, and the analysis of equivalent models defined by the reaction coordinate mapping [34, 35, 36, 37, 38, 39], the chain mapping [40, 41, 42, 43] and the polaron transformation [44, 45, 46, 47, 48, 49, 50]. In alternative, it is also possible construct effective models involving auxiliary degrees of freedom, e.g., the auxiliary density matrices in the hierarchical equations of motion (HEOM) [51, 52, 53, 54, 55, 56, 57, 58, 59, 60], dissipatons [61, 62, 63, 64], and pseudomodes [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78].
@@
In a previous work [79], based on the pseudomode theory [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78], some of us introduced a new class of ancillary damping modes and detailed a rigorous scheme to construct effective models for the numerical simulation of the reduced dynamics of a system and the input-output properties of its environment. These ancillary modes were named “purified pseudomodes” because their states remain pure during the dynamics, in stark contrast to pseudomodes which require a mixed-state description. This method is numerically exact and valid for thermal bosonic environments and for generic linear system-bath interactions, and only requires the spectral decompositions of the bath two-time correlation matrix.
@@
Here, we derive an extended purified pseudomode model for the simulation of open quantum systems in the presence of general nonlinear system-bath interactions, where the bath coupling operator is an analytical function of a linear coupling operator $X$, see the definition in Eq. (2). This topic remains largely unexplored in the literature, despite the notable results in Refs. [80, 81] where the specific case of a system-bath interaction written as the sum of a linear term and a quadratic correction was considered in the context of the HEOM. The analysis of more general nonlinear system-bath interactions, including the specific ones employed in Refs. [80, 81], is more challenging than the linear case because of the additional terms emerging in the Wick factorizations of multi-time bath correlation functions which depend on non-trivial equal-time contractions. In this article, we show that the technical difficulties involved in re-summing all these extra Wick contractions due to nonlinearities are bypassed using the pseudomode theory. In fact, by its very definition, this method offers a natural one-to-one correspondence of multi-time correlation functions between the original bath and the corresponding pseudomode model, allowing the latter to encode any nonlinear statistics of the original bath by mimicking the statistical properties of the linear operator $X$. Despite its generality, the resulting algorithm is easy implementable for the numerical simulation of open quantum systems with non-linear couplings. We numerically illustrate the validity of this method to model the spontaneous decay of a two-level atom strongly coupled to a single-mode lossy cavity. We introduce further potential applications of our method to more complex cases by studying the effects of nonlinearly-coupled phonons on the fluorescence spectrum of a two-level quantum dot.
@@
where $H_{S}$ is the system Hamiltonian and $H_{B}=\sum_{k}\omega_{k}b_{k}^{\dagger}b_{k}$ is the bath Hamiltonian with $b_{k}$ ($b_{k}^{\dagger}$) denoting the annihilation (creation) operator of the bath harmonic mode having frequency $\omega_{k}$, satisfying $b_{k},b_{k^{\prime}}^{\dagger}]=\delta_{kk^{\prime}}$. We consider the following general form for the system-bath interaction
@@
where $\rho_{S}^{I}(t)=e^{iH_{0}t}\rho_{S}(t)e^{-iH_{0}t}$ and $H_{\text{int}}^{I}(t)=e^{iH_{0}t}H_{\text{int}}e^{-iH_{0}t}$ denote the corresponding reduced density operator and system-bath interaction Hamiltonian in the interaction picture, respectively. The formal solution of Eq. (3) can be written [73] as the Dyson series
@@
$\begin{array}[]{lll}\rho_{m}^{I}(t)&\displaystyle=\int_{0}^{t}dt_{m}\cdots\int_{0}^{t}dt_{1}\sum_{i_{1},\cdots,i_{m}\in\{l,r\}}\langle\mathcal{Q}_{t_{m}}^{i_{m}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}\\ &\displaystyle~~~\times\mathcal{T}_{S}\mathcal{S}_{t_{n}}^{i_{n}}\cdots\mathcal{S}_{t_{1}}^{i_{1}}\rho_{S}(0),\end{array}$ (5)
with $\langle\mathcal{Q}_{t_{m}}^{i_{m}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}\equiv\text{Tr}_{B}\Big[\mathcal{T}_{B}\mathcal{Q}_{t_{m}}^{i_{m}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rho_{\text{th}}\Big]$ denoting the multi-time bath correlation functions and the indexes $i_{1},\cdots,i_{m}\in\{l,r\}$ labeling the left- or right-action of the superoperators $\mathcal{S}_{i_{j}}^{t_{j}}$ and $\mathcal{Q}_{t_{j}}^{i_{j}}$. In other words, they are defined as $\mathcal{S}_{t_{j}}^{l}[\cdot]=s(t_{j})[\cdot]$, $\mathcal{S}_{t_{j}}^{r}[\cdot]=-[\cdot]s(t_{j})$, $\mathcal{Q}_{t_{j}}^{l}[\cdot]=Q(t_{j})[\cdot]$, and $\mathcal{Q}_{t_{j}}^{r}[\cdot]=[\cdot]Q(t_{j})$ where $s(t_{j})=\exp(iH_{S}t_{j})s\exp(-iH_{S}t_{j})$ and $Q(t_{j})=\exp(iH_{B}t_{j})Q\exp(-iH_{B}t_{j})$ correspond to the operators $s$ and $Q$ in the interaction picture, respectively. Note that, the minus sign in the definition of $\mathcal{S}_{t_{j}}^{r}$ is necessary to recover the commutators involving $H_{\text{int}}^{I}(t)$ in the Dyson series in Eq. (4). The time-ordering operators $\mathcal{T}_{S}$ and $\mathcal{T}_{B}$ respectively ensure that the superoperators $\mathcal{S}_{i_{j}}^{t_{j}}$ and $\mathcal{Q}_{t_{j}}^{i_{j}}$ are in chronological order.
$\begin{array}[]{lll}\displaystyle C(t)&\displaystyle\equiv\text{Tr}_{B}\left[X(t_{2})X(t_{1})\rho_{\text{th}}\right]\\ &\displaystyle=\int_{0}^{\infty}d\omega\frac{J(\omega)}{\pi}\left[\coth(\beta\omega/2)\cos(\omega t)-i\sin(\omega t)\right],\end{array}$ (6)
where $t=t_{2}-t_{1}$ denote the time difference and $J(\omega)\equiv\pi\sum_{k}\lvert g_{k}\rvert^{2}\delta(\omega-\omega_{k})$ is the spectral density. The time-translational invariance of $C(t)$ is a consequence of the stationarity of the bath initial state $\rho_{\text{th}}$ under the bath free dynamics, manifested in the commutation relation $H_{B},\rho_{\text{th}}]=0$. In addition, $C(t)$ satisfies the time-reversal property $C(t)=C^{*}(-t)$ with the symbol * representing the complex conjugation. Ultimately, the formal series in Eq. (4) can be re-summed to yield a compact expression $\rho_{S}^{I}(t)=\exp(-\mathcal{F}(t,s,C(t))\rho_{S}(0)$ where the influence superoperator $\mathcal{F}(t,s,C(t))$ is explicitly written as [73]
@@
While no compact expression can be found for $\rho_{S}^{I}(t)$ in the general case of nonlinear system-bath interactions considered in Eq. (2), Wick’s theorem [73] still allows to conclude that each multi-time correlation function $\langle\mathcal{Q}_{t_{n}}^{i_{n}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}$ in Eq. (5) can be written as a sum over the product of contractions involving the linear operator $X$, i. e., in terms of the bath correlation function $C(t)$. In the nonlinear case, the resulting factorizations of $\langle\mathcal{Q}_{t_{n}}^{i_{n}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}$ also acquire explicit dependencies on the second moment of $X$, i.e.
@@
As an example, in the quadratic case $Q=X^{2}$, the factorization of $\langle\mathcal{Q}_{t_{2}}^{i_{2}}\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}$ reads
@@
where $\mathcal{X}_{t_{n}}^{l}[\cdot]=X({t_{n}})[\cdot]$, $\mathcal{X}_{t_{n}}^{r}[\cdot]=[\cdot]X({t_{n}})$ with $n=1,2$. The moment $\sigma^{2}$ and the two-time correlation function $C(t)$ appear in the first and second lines, respectively. This explicit dependence on the moment $\sigma^{2}$ in the Wick factorizations of $\langle\mathcal{Q}_{t_{n}}^{i_{n}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}$, has critical and profound consequences on the construction of the purified pseudomode model, constituting a major challenge which we overcome in the next section.
@@ III Purified pseudomode model for nonlinear system-bath interactions
In this section, we explicitly construct a purified pseudomode model that allows to reproduce the reduced system dynamics of the nonlinear open-system model described in Sec. II. To this end, we first present a brief summary of the conventional pseudomode theory in Sec. III.1. We then establish an explicit pseudomode model that reproduces both the two-time correlation in Eq. (6) and the second moment in Eq. (8) in Sec. III.2, based on spectral decompositions of the two-time correlation in Eq. (6). In Sec. III.3, we generalize the purification technique presented in our previous work [79] to reach a purified pseudomode model for the special case $Q=X^{n}$ and furthermore, for the general nonlinear system-bath interaction (2).
@@ III.1 The pseudomode theory
[…]
@@
where $C_{\text{pm}}(t)\equiv\text{Tr}_{\text{pm}}[X_{\text{pm}}(t)X_{\text{pm}}(0)\rho_{\text{vac}}]$ can be solved analytically as [73]
@@
The above consideration can be easily generalized to nonlinear system-bath interactions (2). This is achieved by simply substituting
$Q_{\text{pm}}=\sum_{n=1}^{\infty}\alpha_{n}X_{\text{pm}}^{n}$ (14)
in Eq. (10). The formal correspondence between $Q_{\text{pm}}$ and the original bath coupling operator $Q(X)$ in Eq. (2) has to be such to establish the correspondence $\langle\mathcal{Q}_{t_{m}}^{i_{m}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}=\langle\mathcal{Q}_{\text{pm},t_{m}}^{i_{m}}\cdots\mathcal{Q}_{\text{pm},t_{1}}^{i_{1}}\rangle_{\text{th}}$ between the multi-time correlation functions of the full and effective models (here, the superoperators $\mathcal{Q}_{\text{pm},t_{m}}^{i_{m}}$ are defined analogously as $\mathcal{Q}_{t_{m}}^{i_{m}}$). Using Wick’s theorem, we can then directly conclude that all these equivalences hold when Eq. (12) is imposed.
It is important to note that the equivalence condition in Eq. (12) is not required to be pointwise, i.e., to be valid for all $t\in\mathbb{R}$. More precisely, it can be allowed to be weakly violated, i.e., such that $C(t)\neq C_{\text{pm}}(t)$ over a countable set of times $\mathcal{S}=\{t\in\mathbb{R}|C(t)\neq C_{\text{pm}}(t)\}$. This freedom arises from the property of Riemann integrals and the fact that $C(t)$ always appears as an integrand in Eq. (5). However, we emphasize that the explicit emergence of the second moment in Eq. (8) in the Wick factorizations of the multi-time correlation function $\langle\mathcal{Q}_{t_{m}}^{i_{m}}\cdots\mathcal{Q}_{t_{1}}^{i_{1}}\rangle_{\text{th}}$ for the nonlinear interaction (2) requires a precise approximation of the correlation at $t=0$, i.e., $(t=0)\notin\mathcal{S}$, in stark constrast to the linear case. In Sec. III.2 we show how the freedom described here facilitates the imposition of this constraint at $t=0$ by adding an extra zero-frequency pseudomode with respect to the linear case.